If A and B are dependent events, which of these conditions must be true?
P(A and B)=P(A) + P(B)
P(A and B)= P(B)
P(A)
P(B|A)=P(B)
P(A|B)=P(A)
P(B|A) does not equal P(B)
I think its the first one.

Answer by richard1234(7193)

Not quite. For example, suppose A and B only occur together, that is, B is true if and only if A is true, and suppose P(A) = 1/2. Then P(B) = 1/2, so P(A and B) = 1/2 but P(A) + P(B) = 1.

The last condition must be true -- if P(B|A) = P(B), then A and B are independent.